FEE mounts for 5,6,7,8 disassembled to access adaptor boards. No sign of debris which could cause short on adaptor boards.
Adaptor boards removed and checked against spares with multimeter. Both show same characteristics. Pins on feedthrough checked using multimeter. No longer see -18.5 KOhm between bias pins, no see 140 kOhm which is similar to bias pin to non bias pin from prevous tests.
Camera on interaction chamber removed and replaced with foil, no change.
Ion gauge (Edinburgh group) turned ON
Decided to move detectors using motors. Movement caused change in bias characteristic of the detector. Initial movement was 50 repetitions in and out over which the bias was increased from 10V to 140V. Over this range the detector exhibited diode like behaviour. The difference between the in and position and out position caused a 0.5 uA difference in leakage current. Once stationary, the current was stabel at 140V but after 20s current increased rapidly and tripped.
Increasing bias again to 30V while stationary (out position on motors) gave 7.3 uA , going to 60V tripped again.
Moving the detectors IN but keeping stationary gave a stable current at 30V of 4.2 uA increasing to 70V tripped again. Turning off the gauge reduced the current at 30V to 2.9 uA and 70V to be be ~15.5 uA. Opening an optical window at 70V saw no change in leakage current.
Possibility thermocouple attached to or around the detector may have come loose within the chamber? Thermocouples outisde of chamber removed from thermocouple flange so they don't interact with outside of chamber.
Detector biased to 60V - 10 uA. Detector position moved in and out 20 repetions, during movement current dropped to 2.85 uA. After motion stopped leakage current remained stable at 2.85 uA. Optical window opened no change. Torch light shone into optical window - increase in leakage current to 3.5 uA. Torch off - return to 2.85 uA.
Motion started again for 50 repetitions to test stability of leakage current. Current varies between 2.84-2.88 uA during movement, settles to 2.85 uA once motion stopped.
FEE holder reassembley. Involved turning off bias, unplugging bias cables and rebuilding assembley. FEE holders reaasembled with FEEs reconnected. Water cables, HDMI, ethernet, power connected. Bias turned off and on during reassembley multiple times to check for changes. Each time bias at 60V remained 2.85uA. Each time lower voltages appear to show diode like response in leakage current output.
Pulser connected and FEEs turned back on. Aida05 appears not producing ADC data rates 8 kHz on statistics page -investigate further. Aida06 rate ~70 kHZ can see pulser. Aida07, 08 nn bias show rates ~300 kHZ possible not enough bias for nn side. Rate for all FEEs yesterday (unbiased) was 200-300 kHZ except for AIDA03 disconnected and AIDA02 which wasn't producing ADC data. Photo of Aida06 spectra attached (log scale for pulser).
FEEs turned OFF. Bias left on 60V - 2.85 uA see if stable overnight, stable for approx 2.5 hours before leaving cryring. Detector moved to OUT position, pin back in.
To do - Grounding, reduce noise, waveforms, investigate AIDA05, how high can we bias? |